# Many-Body Localization in the Age of Classical Computing

### v2

@@ VII.3.4 Operator growth
Investigating the Krylov complexity is a possible approach to quantifying the operator spreading. Consider an autocorrelation function $C(t)=\frac{1}{\mathcal{N}}\mathrm{tr}\left(A(t)A\right)$, where $A$ is a local observable, and, for simplicity, we have considered the infinite temperature ensemble of states. Defining a Liouvillian superoperator $\mathcal{L}(\hat{A})\equiv[\hat{H},\hat{A}]$ and using the Baker-Campbell-Hausdorff formula, the autocorrelation function can be written as
$C(t)=\braket{A,\exp(i\mathcal{L}t)A}.$ (57)
where $\braket{A,B}\equiv\frac{1}{\cal N}\mathrm{tr}(A^{\dagger}B)$ is the Hilbert-Schmidt product. The latter equation indicates that the spreading of the operator $A$ under the Hamiltonian dynamics can be analyzed by examination of the sequence of operators $\{\mathcal{L}^{n}A\}$ obtained by consecutive actions of the Liouvillian on $A$. Employing the Lanczos algorithm [353] to construct the Krylov basis of the subspace $\{\mathcal{L}^{n}A\}$ of the operator space, one obtains a sequence of positive numbers, Lanczos coefficients, $\{b_{n}\}$. According to the hypothesis conjectured in [473], in $d$-dimensional chaotic systems, the growth of the Lanczos coefficients is as fast as possible, i.e., linear (up to logarithmic corrections for $d=1$),
@@
An alternative approach introduced in the context of quantum chaos in [38] is to consider the time evolution of observable $A$ in Euclidean time $\tau$, $A(\tau)=e^{\tau H}Ae^{-\tau H}=\exp(\mathcal{L}\tau)A$. In contrast to the real-time evolution, the norm of $A(\tau)$ is not conserved during the Euclidean time evolution, and its growth is related to the class of the dynamics generated by the Hamiltonian $H$. To analyze the growth of $||A(\tau)||$, it suffices to focus on the norms of the operators $\mathcal{L}^{n}A$ generated during the Euclidean time evolution. The following bound was introduced for the dynamics under local interacting Hamiltonian $H$ in dimension $d=1$ [38]
$||\mathcal{L}^{n}A||\leq||A||B_{n}(2)(2J)^{n},$ (59)
@@
where $J$ characterizes the local couplings of $H$, and $B_{n}(2)$ are Bell polynomials. For large $n$, the Bell polynomials scale like $B_{n}(2)\propto(n/\ln(n))^{n}$. For chaotic dynamics, the bound (59) may be expected to be tight, i.e., for ergodic systems the growth of $||\mathcal{L}^{n}A||$ is expected to be faster than exponential and nearly factorial in $n$. Defining a localized interacting system as a system in which the operator $\mathcal{L}^{n}A$ remains exponentially localized with a characteristic length scale $\xi_{\rm loc}$ (equivalently, the operator $A$ remains localized when commuted $n$ times with Hamiltonian), [703] derived the following bound on the norm
$||\mathcal{L}^{n}A||\leq||A||\frac{2^{\xi_{\rm loc}}}{(\xi_{\rm loc}-1)!}(2\xi_{\rm loc}CJ)^{n}.$ (60)
@@
The latter inequality shows that the growth of the norm $||\mathcal{L}^{n}A||$ is at most exponential with $n$ in a localized interacting system. Similarly, for free-fermionic systems, the growth $||\mathcal{L}^{n}A||$ is also at most exponential in $n$. Performing symbolic calculations of the nested commutator $||\mathcal{L}^{n}A||$ for the disordered XXZ spin chain, [703] find that the ratio $\frac{||\mathcal{L}^{n+1}A||}{||\mathcal{L}^{n}A||}$ keeps increasing with $n$, even at disorder strength $W=25$. The largest considered order $n$ is equal to $16$, which corresponds to operators that spread over $L=33$ lattice sites, which allows [703] to conclude that there are no convincing signatures of MBL phase visible at the probed length scale. A comparison of the ratio $\frac{||\mathcal{L}^{n+1}A||}{||\mathcal{L}^{n}A||}$ between interacting and non-interacting cases shows that the growth is similar in the two cases at small $n\lesssim 8$, and it then increases its pace significantly in the interacting model. This finding parallels the observations that the non-interacting limit of the disordered XXZ spin chains provides a good proxy for the dynamics on relatively short time scales in the interacting case, provided that the disorder is sufficiently strong.
@@
Finding quantitative relations of the Euclidean time operator spreading with the persistent decays of local autocorrelation functions observed in the time evolution approaches is an interesting avenue for further research. Understanding the properties of operator spreading in models which are by definition MBL, such as [131], or obtaining mathematical insights into operator spreading in disordered systems [192, 193], could shed new light on distinguishing the regime of slow dynamics in strongly disordered spin chains from the properties of the dynamics in MBL phase in interacting models.
@@ VII.3.5 Dynamics of survival probabilities
Expressing the initial state |$\ket{\psi(0)}=\sum_{n}c_{n}\ket{n}$ in the eigenbasis {|$\{\ket{n}\}$ of the Hamiltonian $H$, and performing an average $\braket{.}$ over disorder realizations, the survival probability can be rewritten as
@@
where the coefficients $|c_{n}|^{2}$ specify the distribution of the initial state in the eigenbasis of the Hamiltonian. The latter expression can be made identical to the definition of the SFF in Eq. (30) when $|c_{n}|^{2}$ are replaced by the Gaussian filter $g(E)$. Consequently, in the ergodic regime, the behavior of the survival probability at long times resembles the ramp-plateau structure of the SFF. Before the increase of $P(t)$ associated with the onset of ramp behavior, the survival probability has a broad minimum called a correlation hole [366]. Pin-pointing the position of the correlation hole allows for an alternative definition of a Thouless time [556] at which the spread of the initial state in the many-body Hilbert space is complete. The Thouless time defined in this fashion was shown to scale exponentially with the system size in the ergodic regime of disordered spin chains, starkly contrasting the $L^{2}$ scaling found in the Thouless time from SFF. Notably, the correlation hole is associated with a "bulge" in the time evolution of the entanglement entropy $S(t)$ described by [664]. Both the correlation hole and the entanglement entropy "bulge" are dynamic manifestations of quantum chaos and are sensitive even to exponentially weak integrability terms [551]. The correlation hole, measurable experimentally with survival probability or correlation functions [152], allows for the detection of the onset of chaotic dynamics in the regime of slow dynamics in disordered spin chains.
Despite considerable activities in studies of survival probabilities, no sharp prediction for the onset of MBL (or its absence) has been formulated so far. In contrast, for the quantum sun model Eq. (11), in which the transition point is well established, [280] showed that the survival probability, when rescaled by its asymptotic values as routinely done in studies of the SFF, exhibits scale invariance at the transition. The latter provides an intriguing perspective to be explored in other interacting models as well.
The survival probability can also be viewed as a return probability on the Fock-space graph. In this case it can be directly linked with properties of correlation functions of local observables [459]. Notably, higher-order correlations on the Fock-space graph carry information about entanglement in the system [531], providing an alternative perspective on the slow growth of entanglement in the disordered spin chains. The transport of probability on the Fock-space graph [142] is an alternative framework for studying universal features of the slow dynamics in the disordered spin chains. For example, the slow sub-diffusive dynamics in strongly disordered spin chains can be interpreted as originating from heavy-tailed distributions of the escape times in the vicinity of the depinning transition in directed polymers in random media [77].
@@ VIII.1 Transport in the ergodic regime
A general expectation for quantum, as well as classical, many-body chaotic systems is that their dynamics are diffusive [545, 412]. The presence of diffusion in many-body classical systems can be linked to the positivity of some Lyapunov exponents, which necessarily arise when integrability is broken. Such arguments, which in a classical system can be made by a combination of analytical and large-system sizes numerics (thousands of microscopic constituents), do not have direct quantum-mechanical counterparts. In quantum, non-integrable systems one has to necessarily resort to smaller system sizes, and investigate transport with almost no support from analytical calculations. Even in these circumstances, however, one might be led to believe that diffusive transport must be the most common behavior of an ergodic quantum system (at small disorder, weak interactions), with sub- and super-diffusion or ballistic transport relegated to a few counterexamples. In this section we review this question in the context of systems exhibiting MBL regime, and we show that the counterexamples are more than a few.
@@
The presence of diffusion in 1D non-integrable lattice models was substantiated by several complementary methodological approaches [64], including, e.g., tensor network investigations [510, 509] as well as time evolution results [324, 307, 624]. Moreover, diffusion is observed when weak interactions are introduced to localized quasiperiodic systems [737, 674]. However, an example of a non-integrable spin ladder, which supports a ballistic magnetization transport in its invariant subspaces and whose transport is otherwise diffusive, was put forward in [696]. Hence, there are examples of chaotic, ergodic systems with spectral statistics following RMT predictions, that support transport types that are different than diffusive.
@@
Integrable systems with a single impurity exhibit spectral properties that adhere to predictions of RMT [548, 50]. However, as shown on the example of the XXZ spin chain with a single impurity, such system hosts a ballistic transport [88]. Investigations of kinetic constraints in 1D many-body systems, such as the $PXP$ model [the paradigmatic model of quantum many-body scars [575]] provide further evidence for transport that is different from diffusive in the ergodic phase. For example, the analysis of energy transport in the kinetically constrained $PXP$ model using extensive tensor network investigations indicates that the transport is superdiffusive, and the dynamical critical exponent is $z\approx 3/2$ [377]. The appearance of a superdiffusive transport in the $PXP$ model is aligned with the view of its proximity to the integrable point [312], and with the claim about the lack of MBL phase in this type of constrained spin chains [594]. Similar findings about anomalous diffusion have been reported for ergodic systems with dipole and higher moment conserving lattice systems [203] as well as in models with sparse interactions [160]. From that perspective, understanding conditions in which non-diffusive transport may emerge in ergodic systems remains an interesting question for further research.
@@ VIII.1.1 Spin conductivity
where $j=\sum_{l}j_{l}$, with $j_{l}$ defined in (67) is the spin current operator [^31] If one considers spinless fermions, $\sigma(\omega)$ corresponds to charge conductivity., and the proportionality factor in Eq. (63) includes the temperature $T$, such that the non-trivial quantity in the high-temperature limit $T\to\infty$ is $\tilde{\sigma}(\omega)=T\sigma(\omega)$; in what follows, we set $\tilde{\sigma}\to\sigma$. The dynamical conductivity $\sigma(\omega)$ is the spectral function (40) of the current operator $j$. However, the matrix elements $\braket{n}{j}{m}$ of the current operator $j$ in the disordered spin chains differ significantly from the matrix elements of $S^{z}_{j}$, leading to the expected small frequency behavior
@@
where $\sigma_{0}$ is the DC conductivity, $\zeta$ is a constant and the exponent $\bar{\alpha}$ can be related to the dynamical critical exponent as $\bar{\alpha}=1-\frac{2}{z}$ [388]. To extract the DC conductivity, one must first consider the limit of $L\to\infty$, followed by $\omega\to 0$. This task is feasible in the ergodic regime, in which the dynamical conductivity $\sigma(\omega)$ at small frequencies behaves in a fashion independent of the broadening $\eta$ of the Dirac delta functions used for numerical evaluation of Eq. (63), see also the discussions in [503, 388]. Analysis of the disordered spin chains [502, 49] finds that the diffusion constant $D_{0}\propto\sigma_{0}$ in the ergodic regime decreases as $D_{0}\propto e^{-aW}$, i.e. exponentially with the disorder strength. Linking $D_{0}$ with a parameter that quantifies the sensitivity of energy levels to "twist" of boundary conditions (see Sec. VI.3.2), Ref. [502] argues that the deviations from the RMT universality in disordered spin chains are observable when the diffusion constant $D_{0}$ starts to be comparable to the average level spacing. This occurs at disorder strength $W\propto L$, similar to $W^{T}(L)\propto L$ extracted in Sec.VI.1, due to the combination of the exponential decrease of $D_{0}$ with $W$, and the exponential decrease of the average level spacing with system size $L$.
@@ VIII.1.2 Unitary time evolution
Early studies of disordered XXZ spin chain dynamics have reported anomalous, sub-diffusive dynamics at disorder strengths below the MBL regime $W\lessapprox W^{*}(L)$ [44]. Similar conclusions have been obtained from analyzing the density autocorrelation function $C(t)$ in [394]. Importantly, the subsequent studies [55, 702] of the two-point correlation function $\Phi(x,t)$ defined in Sec. VII.1.3 indicate, instead, a prominent decrease of the extracted dynamical exponent $z$ with system size $L$. For instance, the maximal value of $z$ (extracted after the initial transient period and before the late time saturation) changes between $z\approx 4.2$ for $L=16$, through $z\approx 3.4$ at $L=24$ down to $z\approx 2.95$ for $L=32$ at $W=1.5$ [702]. While at $L=32$ the dynamical exponent $z$ is still markedly larger from the diffusive limit $z=2$, the significance of its drift with system size prevents one from excluding a diffusive scenario at $W=1.5$. With increasing the disorder strength, the system size drift of $z$ becomes weaker but is still observable. For instance at $W=3$, Ref. [702] finds $z\approx 6.4$ at $L=16$ and $z\approx 5.2$ at $L=28$. While the slowdown of the dynamics is evident, the available numerical results for $\Phi(x,t)$ disallow clear conclusions about the class of the ergodic dynamics at $W=3$ in the disordered XXZ spin chain [^32] Similarly, at $W=3$, there is no stable power-law exponent $\beta$ governing the decay of $C(t)$. Instead, a monotonic increase of $\beta$ with system size $L$ is observed; see Fig. 10(b)..
Similar difficulties are faced even in transport studies in non-interacting systems. For example, let us consider Anderson model (5) on a 3D regular (cubic) lattice with on-site energies $\epsilon_{i}$ being random, independent variables uniformly distributed in the interval $[-W/2,W/2]$. It is well established that this model has Anderson transition at $W\approx 16.5$ [611] and that the dynamics are subdiffusive at the transition, with the dynamical exponent $z=3$ [452].
Due to the absence of interactions in the Anderson model, we consider a quantity related to $\Phi(x,t)$ by probing the time evolution of an initial state |$\ket{\psi_{0}}$ with a single particle completely localized at a given lattice site. Calculating the time evolved state $|\psi_{0}(t)\rangle=e^{-i\hat{H}t}|\psi_{0}\rangle$, we compute the mean square displacement
@@
Figure 13: Subdiffusion at Anderson transition in $3D$ (left) and $5D$ (right) systems. Time dependent $\alpha(t)$ function for 3D (left) and 5D (right) Anderson models for various disorder strengths $W$. In the 3D case, results for the system size $L=80,120,160,240$ are denoted by progressively thicker lines, whereas in the 5D case thin (thick) lines correspond to $L=20$ ($L=30$). For $t\to\infty$ the flowing exponent vanishes, $\alpha(t)\to 0$, when the wave-packet finishes spreading over the entire lattice. Reprinted from [593].
@@
The results for Anderson models indicate that the time and length scales required to assess the character of the dynamics correctly may be substantial. The same conclusion applies also to disordered classical Heisenberg chains  [409]. Hence, from this perspective in it not surprising to observe the finite size drifts of the exponent $z$ towards $z=2$, such as those found in the ergodic regime of the disordered Heisenberg spin chain [55, 702], and their interpretation must be aligned with results from other transport measures obtained from complementary methods.

### v3

@@ VII.3.4 Operator growth
Investigating the Krylov complexity is a possible approach to quantifying the operator spreading. Consider an autocorrelation function $C(t)=\frac{1}{\mathcal{N}}\mathrm{tr}\left(\hat{A}(t)\hat{A}\right)$, where $A$ is a local observable, and, for simplicity, we have considered the infinite temperature ensemble of states. Defining a Liouvillian superoperator $\mathcal{L}(\hat{A})\equiv[\hat{H},\hat{A}]$ and using the Baker-Campbell-Hausdorff formula, the autocorrelation function can be written as
$C(t)=\braket{\hat{A},\exp(i\mathcal{L}t)\hat{A}};,$ (57)
where $\braket{\hat{A},\hat{B}}\equiv\frac{1}{\cal N}\mathrm{tr}(\hat{A}^{\dagger}\hat{B})$ is the Hilbert-Schmidt inner product. The latter equation indicates that the spreading of the operator $A$ under the Hamiltonian dynamics can be analyzed by examination of the sequence of operators $\{\mathcal{L}^{n}\hat{A}\}$ obtained by consecutive actions of the Liouvillian on $\hat{A}$. Employing the Lanczos algorithm [382] to construct the Krylov basis of the subspace $\{\mathcal{L}^{n}\hat{A}\}$ of the operator space, one obtains a sequence of positive numbers, Lanczos coefficients, $\{b_{n}\}$. According to the hypothesis conjectured in [511], in $d$-dimensional chaotic systems, the growth of the Lanczos coefficients is as fast as possible, i.e., linear (up to logarithmic corrections for $d=1$),
@@
An alternative approach introduced in the context of quantum chaos in [40] is to consider the time evolution of observable $A$ in Euclidean time $\tau$, $\hat{A}(\tau)=e^{\tau\hat{H}}\hat{A}e^{-\tau\hat{H}}=\exp(\mathcal{L}\tau)\hat{A}$. In contrast to the real-time evolution, the norm of $\hat{A}(\tau)$ is not conserved during the Euclidean time evolution, and its growth is related to the class of the dynamics generated by the Hamiltonian $\hat{H}$. To analyze the growth of $||\hat{A}(\tau)||$, it suffices to focus on the norms of the operators $\mathcal{L}^{n}A$ generated during the Euclidean time evolution. The following bound was introduced for the dynamics under local interacting Hamiltonian $\hat{H}$ in dimension $d=1$ [40],
$||\mathcal{L}^{n}\hat{A}||\leq||\hat{A}||B_{n}(2)(2J)^{n},$ (59)
@@
where $J$ characterizes the local couplings of $\hat{H}$, and $B_{n}(2)$ are Bell polynomials. For large $n$, the Bell polynomials scale like $B_{n}(2)\propto(n/\ln(n))^{n}$. For chaotic dynamics, the bound (59) may be expected to be tight, i.e., for ergodic systems the growth of $||\mathcal{L}^{n}\hat{A}||$ is expected to be faster than exponential and nearly factorial in $n$. Defining a localized interacting system as a system in which the operator $\mathcal{L}^{n}\hat{A}$ remains exponentially localized with a characteristic length scale $\xi_{\rm loc}$ (equivalently, the operator $\hat{A}$ remains localized when commuted $n$ times with Hamiltonian), [759] derived the following bound on the norm
$||\mathcal{L}^{n}\hat{A}||\leq||\hat{A}||\frac{2^{\xi_{\rm loc}}}{(\xi_{\rm loc}-1)!}(2\xi_{\rm loc}CJ)^{n}.$ (60)
@@
The latter inequality shows that the growth of the norm $||\mathcal{L}^{n}\hat{A}||$ is at most exponential with $n$ in a localized interacting system. Similarly, for free-fermionic systems, the growth $||\mathcal{L}^{n}\hat{A}||$ is also at most exponential in $n$. Performing symbolic calculations of the nested commutator $||\mathcal{L}^{n}\hat{A}||$ for the disordered XXZ spin chain, [759] find that the ratio $\frac{||\mathcal{L}^{n+1}\hat{A}||}{||\mathcal{L}^{n}\hat{A}||}$ keeps increasing with $n$, even at disorder strength $W=25$. The largest considered order $n$ is equal to $16$, which corresponds to operators that spread over $L=33$ lattice sites, which allows [759] to conclude that there are no convincing signatures of the MBL phase visible at the probed length scale. A comparison of the ratio $\frac{||\mathcal{L}^{n+1}\hat{A}||}{||\mathcal{L}^{n}\hat{A}||}$ between interacting and non-interacting cases shows that the growth is similar in the two cases at small $n\lesssim 8$, and it then increases its pace significantly in the interacting model. This finding parallels the observations that the non-interacting limit of the disordered XXZ spin chains provides a good proxy for the dynamics on relatively short time scales in the interacting case, provided that the disorder is sufficiently strong.
@@
Finding quantitative relations of the Euclidean time operator spreading with the persistent decays of local autocorrelation functions observed in the time evolution approaches is an interesting avenue for further research. Understanding the properties of operator spreading in models which are by definition MBL, such as [136], or obtaining mathematical insights into operator spreading in disordered systems [202, 203]404040 Arguments for the stability of slow Hamiltonian dynamics to generic local perturbations are presented in [714]., could shed new light on distinguishing the regime of slow dynamics in strongly disordered spin chains from the properties of the dynamics in MBL phase in interacting models.
@@ VII.3.5 Dynamics of survival probabilities
Expressing the initial state |$\ket{\psi(0)}=\sum_{n}c_{n}\ket{n}$ in the eigenbasis {|$\{\ket{n}\}$ of the Hamiltonian $\hat{H}$, and performing an average $\braket{\cdots}$ over disorder realizations, the survival probability can be rewritten as
@@
where the coefficients $|c_{n}|^{2}$ specify the distribution of the initial state in the eigenbasis of the Hamiltonian. The latter expression can be made identical to the definition of the SFF in Eq. (30) when $|c_{n}|^{2}$ are replaced by the Gaussian filter $g(E)$. Consequently, in the ergodic regime, the behavior of the survival probability at long times resembles the ramp-plateau structure of the SFF. Before the increase of $P(t)$ occurs (the increase is associated with the onset of ramp behavior), the survival probability has a broad minimum called a correlation hole [399]. Pin-pointing the position of the correlation hole allows for an alternative definition of a Thouless time [603], at which the spread of the initial state in the many-body Hilbert space is complete. The Thouless time defined in this fashion was shown to scale exponentially with the system size in the ergodic regime of disordered spin chains, starkly contrasting the $L^{2}$ scaling found in the Thouless time from SFF. Notably, the correlation hole is associated with a "bulge" in the time evolution of the entanglement entropy $S(t)$ described by [717]. Both the correlation hole and the entanglement entropy "bulge" are dynamic manifestations of quantum chaos and are sensitive even to exponentially weak integrability terms [597]. The correlation hole, measurable experimentally via survival probability or correlation functions [158], allows for the detection of the onset of chaotic dynamics in the regime of slow dynamics in the disordered spin chains.
Despite considerable activities in studies of survival probabilities, no sharp prediction for the onset of MBL (or its absence) has been formulated so far. In contrast, for the quantum sun model in Eq. (11), in which the transition point is well established, [299] showed that the survival probability, when rescaled by its asymptotic values as routinely done in studies of the SFF, exhibits scale invariance at the transition. The latter provides an intriguing perspective to be explored in other interacting models as well.
The survival probability can also be viewed as a return probability on the Fock-space graph. In this case it can be directly linked with the properties of correlation functions of local observables [497]. Notably, higher-order correlations on the Fock-space graph carry information about entanglement in the system [577], providing an alternative perspective on the slow growth of entanglement in the disordered spin chains. The transport of probability on the Fock-space graph [147] is an alternative framework for studying universal features of the slow dynamics in the disordered spin chains. For example, the slow sub-diffusive dynamics in strongly disordered spin chains can be interpreted as originating from heavy-tailed distributions of the escape times in the vicinity of the depinning transition, found in directed polymers in random media [80].
@@ VIII.1 Transport in the ergodic regime
A general expectation for quantum, as well as classical, many-body chaotic systems is that their dynamics are diffusive [591, 448]. The presence of diffusion in many-body classical systems can be linked to the positivity of some Lyapunov exponents, which necessarily arise when integrability is broken. Such arguments, which in a classical system can be made by a combination of analytical and large system size numerics (thousands of microscopic constituents), do not have direct quantum-mechanical counterparts. In quantum non-integrable systems, one has to necessarily resort to smaller system sizes, and investigate transport with almost no support from analytical calculations. Even in these circumstances, however, one might be led to believe that diffusive transport must be the most common behavior of an ergodic quantum system, with sub- and super-diffusion or ballistic transport relegated to a few counterexamples. In this Section we review this question in the context of systems exhibiting MBL regime, and we show that the counterexamples are more than a few.
@@
The presence of diffusion in 1D non-integrable lattice models was substantiated by several complementary methodological approaches [67], including, e.g., tensor network investigations [551, 550] as well as time evolution results [350, 333, 674]. Moreover, diffusion is observed when weak interactions are introduced to localized quasiperiodic systems [794, 728]. However, an example of a non-integrable spin ladder, which supports a ballistic magnetization transport in its invariant subspaces and whose transport is otherwise diffusive, was put forward in [752]. Hence, there are examples of chaotic, ergodic systems with spectral statistics following RMT predictions, which support transport types that are different from diffusive.
@@
Integrable systems with a single impurity exhibit spectral properties that adhere to predictions of RMT [594, 53]. However, as shown on the example of the XXZ spin chain with a single impurity, such system hosts a ballistic transport [91]. Investigations of kinetic constraints in 1D many-body systems, such as the $PXP$ model [the paradigmatic model of quantum many-body scars [622]] provide further example of transport that is different from diffusive in the ergodic phase. For example, the analysis of energy transport in the kinetically constrained $PXP$ model, using extensive tensor network investigations, indicates that the transport may be superdiffusive, and the dynamical critical exponent is $z\approx 3/2$ [413]. The appearance of a superdiffusive transport in the $PXP$ model is aligned with the view of its proximity to the integrable point [338], and with the claim about the lack of MBL phase in this type of constrained spin chains [643]. Similar findings about anomalous diffusion have been reported for ergodic systems with dipole and higher moment conserving lattice systems [215], in models with sparse interactions [167] and in kinetically constrained quantum circuits [658]. From that perspective, understanding conditions in which non-diffusive transport may emerge in ergodic systems remains an interesting question for further research.
@@ VIII.1.1 Spin conductivity
where $\hat{j}=\sum_{l}\hat{j}_{l}$, with $j_{l}$ defined in (67) is the spin current operator [^41] If one considers spinless fermions, $\sigma(\omega)$ corresponds to charge conductivity., and the proportionality factor in Eq. (63) includes the temperature $T$, such that the non-trivial quantity in the high-temperature limit $T\to\infty$ is $\tilde{\sigma}(\omega)=T\sigma(\omega)$; in what follows, we set $\tilde{\sigma}\to\sigma$. The dynamical conductivity $\sigma(\omega)$ is the spectral function (40) of the current operator $\hat{j}$. However, the matrix elements $\braket{n}{\hat{j}}{m}$ of the current operator $\hat{j}$ in the disordered spin chains differ significantly from the matrix elements of $\hat{S}^{z}_{j}$, leading to the expected small frequency behavior
@@
where $\sigma_{0}$ is the DC conductivity, $\zeta$ is a constant and the exponent $\bar{\alpha}$ can be related to the dynamical critical exponent as $\bar{\alpha}=1-\frac{2}{z}$ [424]. To extract the DC conductivity, one must first consider the limit of $L\to\infty$, followed by $\omega\to 0$. This task is feasible in the ergodic regime, in which the dynamical conductivity $\sigma(\omega)$ at small frequencies behaves in a fashion independent of the broadening $\eta$ of the Dirac delta functions used for numerical evaluation of Eq. (63), see also the discussions in [544, 424]. Analysis of the disordered spin chains [543, 52] finds that the diffusion constant $D_{0}$ (satisfying $D_{0}\propto\sigma_{0}$ due to Einstein relation) in the ergodic regime decreases as $D_{0}\propto e^{-aW}$, i.e., exponentially with the disorder strength. Linking $D_{0}$ with a parameter that quantifies the sensitivity of energy levels to "twist" of boundary conditions (see Sec. VI.3.2), Ref. [543] argued that the deviations from the RMT universality in disordered spin chains are observable when the diffusion constant $D_{0}$ starts to be comparable to the average level spacing. This occurs at disorder strength $W\propto L$, similar to $W^{T}(L)\propto L$ extracted in Sec.VI.1, due to the combination of the exponential decrease of $D_{0}$ with $W$, and the exponential decrease of the average level spacing with system size $L$.
@@ VIII.1.2 Unitary time evolution
Early studies of the dynamics in the disordered XXZ spin chain reported anomalous, sub-diffusive dynamics at disorder strengths below the MBL regime, $W\lessapprox W^{*}(L)$ [47]. Similar conclusions have been obtained from analyzing the density autocorrelation function $C(t)$ in [431]. Importantly, the subsequent studies [58, 758] of the two-point correlation function $\Phi(x,t)$ defined in Sec. VII.1.3 indicate, instead, a prominent decrease of the extracted dynamical exponent $z$ with system size $L$. Denoting the $n$-th moment of $\Phi(x,t)$ by $\braket{x^{n}}=\sum_{x}x^{n}\,\Phi(x,t)$, the density spreading may be quantified by $\Delta x(t)=(\braket{x^{2}}-\braket{x}^{2})^{1/2}\propto t^{1/z}$. For instance, the maximal value of $z$ (extracted after the initial transient period and before the late time saturation) changes between $z\approx 4.2$ for $L=16$, through $z\approx 3.4$ at $L=24$ down to $z\approx 2.95$ for $L=32$ at $W=1.5$ [758]. While at $L=32$ the dynamical exponent $z$ is still markedly larger from the diffusive limit $z=2$, the significance of its drift with system size prevents one from excluding a diffusive scenario at $W=1.5$. With increasing the disorder strength, the system size drift of $z$ becomes weaker but is still observable. For instance at $W=3$, [758] found $z\approx 6.4$ at $L=16$ and $z\approx 5.2$ at $L=28$. While the slowdown of the dynamics is evident, the available numerical results for $\Phi(x,t)$ disallow clear conclusions about the class of the ergodic dynamics at $W=3$ in the disordered XXZ spin chain [^42] Similarly, at $W=3$, there is no stable power-law exponent $\beta$ governing the decay of $C(t)$. Instead, a monotonic increase of $\beta$ with system size $L$ is observed; see Fig. 10(b)..
Similar difficulties are faced even in transport studies in non-interacting systems. For example, let us consider Anderson model (5) on a 3D regular (cubic) lattice with on-site energies $\epsilon_{i}$ being random, independent variables uniformly distributed in the interval $[-W/2,W/2]$. It is well established that this model has the Anderson transition at $W\approx 16.5$ [661] and that the dynamics are subdiffusive at the transition, with the dynamical exponent $z=3$ [490].
Due to the absence of interactions in the Anderson model, we consider a quantity related to $\Phi(x,t)$ by probing the time evolution of an initial state |$\ket{\psi_{0}}$ with a single particle completely localized at a given lattice site. Calculating the time evolved state $|\psi_{0}(t)\rangle=e^{-i\hat{H}t}|\psi_{0}\rangle$, we compute the mean square displacement,
@@
Figure 13: Subdiffusion at Anderson transition in $3$D (left) and $5$D (right) hypercubic lattices. Results show the time dependent $\alpha(t)$ function for 3D (left) and 5D (right) Anderson models for various disorder strengths $W$. In the 3D case, results for the system size $L=80,120,160,240$ are denoted by progressively thicker lines, whereas in the 5D case thin (thick) lines correspond to $L=20$ ($L=30$). For $t\to\infty$ the flowing exponent vanishes, $\alpha(t)\to 0$, when the wave-packet finishes spreading over the entire lattice. Reprinted from [642].
@@
The results for Anderson models indicate that the time and length scales required to assess the character of the dynamics correctly may be substantial. The same conclusion applies also to disordered classical Heisenberg chains  [445]. Hence, from this perspective it is not surprising to observe the finite size drifts of the exponent $z$ towards $z=2$, such as those found in the ergodic regime of the disordered XXZ spin chain [58, 758], and their interpretation must be aligned with results from other transport measures obtained from complementary methods.